$SPAD/src/input synonym2.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- This file contains the standard system defined system command
-- synonyms for the Scratchpad II system. It is normally read into
-- the system by adding the line
--         )read synonym input )ifthere )quiet
-- to the start-up profile SPADPROF INPUT (on CMS) or spadprof.input
-- (on AIX).
-- If you wish to have a private list of synonyms, it is suggested
-- that you create a file MYSYNS INPUT (on CMS) or mysyns.input (on AIX).
-- This will also automatically be read into the system by the start-up
-- profile by the line
--         )read mysyns input )ifthere )quiet
)synonym  ?            what commands
)synonym  ap           what things
)synonym  apr          what things
)synonym  apropos      what things
--)synonym  bug          system spadnote
)synonym  cache        set functions cache
)synonym  cl           clear
)synonym  cls          zsystemdevelopment )cls
)synonym  cms          system
--)synonym  CMS          system
--)synonym  cp           system cp
--)synonym  CP           system cp
)synonym  co           compiler
)synonym  d            display
)synonym  dep          display dependents
)synonym  dependents   display dependents
)synonym  disc         system cp disc
)synonym  e            edit
--)synonym  exec         system exec
)synonym  fc           zsystemdevelopment )c
)synonym  fct          zsystemdevelopment )ct
)synonym  fe           zsystemdevelopment )e
)synonym  fec          zsystemdevelopment )ec
)synonym  fect         zsystemdevelopment )ect
)synonym  fns          exec spadfn
)synonym  fortran      set output fortran
--)synonym  glos         system exec getgloss
)synonym  h            help
)synonym  ht           system hypertex &
)synonym  include      compile
)synonym  kclam        boot clearClams ( )
)synonym  killcaches   boot clearConstructorAndLisplibCaches ( )
--)synonym  logoff       system cp logoff
--)synonym  note         system note

)synonym  patch        zsystemdevelopment )patch
)synonym  pause        zsystemdevelopment )pause
--)synonym  peek         system conpeek
)synonym  prompt       set message prompt
--)synonym  rdr          system checkrdr
)synonym  recurrence   set functions recurrence
)synonym  restore      history )restore
)synonym  save         history )save
)synonym  seq          set streams
)synonym  sequence     set streams
)synonym  storage      set message storage
)synonym  str          set streams
)synonym  streams      set streams
--)synonym  tell         system tell
)synonym  time         set message time
)synonym  type         set message type
)synonym  up           zsystemdevelopment )update
)synonym  version      zsystemdevelopment )version
)synonym  w            what
)synonym  wc           what categories
)synonym  wd           what domains
)synonym  wp           what packages
)synonym  ws           what synonyms
--)synonym  x            system editarb

———-

References^{[1] nothing}

5